OpenStudy (anonymous):

All of the following are equivalent except _____. -5(x- 1) -5x+ 5 5 - 5x (5 - 5)x

OpenStudy (anonymous):

@trumpetmaster7777

OpenStudy (ivettef365):

Well solve all of them -5(x-1) = -5x + 5 -5x+5 = -5x + 5 5 - 5x = -5x +5 (5-5)x = -5x + 5x